i'm trying to solve this problem but i really have no idea how to do it.
How many milliliters of a 6% alcohol solution must be added to 12 milliliters of a 31% alcholol solution to produce a 16% alcohol solution?

Answer by Earlsdon(6294)

Let x = the required number of milliliters of 6% alcohol solution.
Let's find out how much alcohol we have.
The number of milliliters of alcohol in the x ml. of the 6% solution is:
x(0.06) and in the 12 ml. of the 31% solution we have 12(0.31) ml. of alcohol.
When added together, these are to equal (12+x)(0.16)ml. of alcohol, so we can write the equation to find x.
0.06x+0.31(12) = 0.16(12+x) Simplify and solve for x.
0.06x+3.72 = 1.92+0.16x Subtract 0.06x from both sides.
3.72 = 1.92+0.1x Subtract 1.92 from both sides.
1.8 = 0.1x Divide by 0.1
x = 18
You will need to add 18ml. of 6% alcohol solution to the 12ml. of the 31% alcohol solution to obtain (12+x = 30)ml. of 16% alcohol solution.